The length of a rectangle is 3m less than twice the width, and the area of the rectangle is 14m^2 . Find the dimensions of the rectangle.
Answer by mananth(16946)

let width be x
length 2x-3
Area = Length * width
x(2x-3) =14
2x^2-3x =14
2x^2-3x-14=0
2x^2-7x+4x-14 =0
x(2x-7) +2(2x-7)=0
(2x-7)(x+2)=0
x=7/2 OR -2
ignore negative
width = 3.5 m
length = 2x-3 = 2*7/2 -3 =4
Length = 4 m